Find the total surface area of a right triangular prism whose base sides are 5cm,12cm and 13cm,and its height is 11cm. the base is a right angled triangle..
I need your help!!!!Help me now plzz!!!!!

Answer by Alan3354(21550)

Find the total surface area of a right triangular prism whose base sides are 5cm,12cm and 13cm,and its height is 11cm. the base is a right angled triangle..
--------------
The base's area is 5*12/2 = 30. The "top" is also 30.
The other 3 faces have areas of:
11*5, 11*12 and 11*13
= 11*(5+12+13) = 11*30
Area = 330 + 30 + 30
Area = 390 sq cm